D=C-s/n (for s)Depreciation
Hints:
Isolate s on one side of the equals sign.
Your answer will be a formula, not a number.

Answer by stanbon(75887)

D=C-s/n
s/n=C-D
s = n(C-D)
========
Cheers,
Stan H.